# Gravitational Waves Induced by Scalar Perturbations with a Broken Power-law Peak

### v1

@@ III.1 Features and extreme cases
$\lambda=\alpha/\beta$ and $K(\kappa)=\lambda\kappa^{\alpha+\beta}$ are introduced for simplicity, and $\delta$ is the relative error of the power expansion to $K^{n}$ term. To the leading order, the series is reduced to $n=0$ term. We assume that the power-law feature is broken when the leading order of the asymptotic expansion reaches the peak value, where $\kappa$ and $\mathcal{P_{R}}(\kappa)$ satisfy:
@@
$\mathcal{P_{R}}(\kappa_{-})>\frac{A}{2},~{}\mathcal{P_{R}}(\kappa_{+})>\frac{A}{2},~{}\mathcal{P_{R}}(\kappa_{+})\mathcal{P_{R}}(\kappa_{-})>\frac{A^{2}}{2},$
where $\kappa_{-}$ and $\kappa_{+}$ are the critical values of IR and UV expansion respectively, and $\kappa_{0}$ is the boundary of the convergence domain of Eq. (15). The relation indicates that even if we connect the leading order tail with the transition part replaced by the peak value, the relative error is smaller than 1, and approaches 1 only if the power index is much larger than that on the other side, i.e., $|\ln(\alpha/\beta)|\gg 1$. In order to elevate the approximation to a higher accuracy, we introduce a variable transformation that maintains $\mathcal{P_{R}}$:
$\mathcal{P_{R}}(\alpha,\beta;\kappa)=\mathcal{P_{R}}(\lambda,1,\kappa^{\beta})=\mathcal{P_{R}}(\lambda^{-1},1,\kappa^{-\alpha}),$ (17)
Therefore, once the approximation of $\mathcal{P_{R}}(\lambda,1,\kappa^{\beta})$ on $10<\kappa<$ 1 and $\lambda>0$ is finished, it can be extended to the complete parameter space $\alpha,\beta,\kappa>0$ with the accuracy preserved. The the selection is based on that $\kappa_{-}^{\beta}=(1+\lambda)^{-1/\lambda}>e^{-1}$, which make it convenient to extend the power-law tail to the peak with a quadratic polynomial. On the region $\kappa<\kappa_{-}$, the deviation of the power-law tail from $\mathcal{P_{R}}$ is roughly to the order $(2\lambda+1)$, and consequently a general $\mathcal{P_{R}}(\kappa)$ can be approximated by:
@@
Additionally, $\mathcal{P_{R}}^{2}(\kappa)$ will be used in the calculation of $\Omega_{\mathrm{GW}}$, since it is associated with the $p=|\bm{k}-\bm{p}|$ slice in Eq. (5), which separates the integral into two identical parts:
@@
We have used that $\kappa_{-}(\lambda,1)=\kappa_{-}^{\beta}(\alpha,\beta)$.
@@
III.2 Analytical approach to SIGW
In this part, we will calculate the GW spectrum derived in Sec. II, with the scalar spectrum $\mathcal{P_{R}}$ having a broken power-law feature as Eq. (14). To make use of the symmetry, we perform the transformation $x=\frac{1}{\sqrt{2}}(u+v),~{}y=\frac{1}{\sqrt{2}}(u-v)$ on the integral in Eq. (12):
where we have defined $q=\kappa x/\sqrt{2}$ for simplicity. Obviously, most of the factors of $F(x,y)$ are not sensitive to $y$ since $y^{2}<1/2$, except for the region near the zero point of $(x^{2}+y^{2}-3)$ and $(x^{2}-y^{2})$. The former contributes little deviation to $\Omega_{\mathrm{GW}}$ since it is suppressed by the integration on the nearby peak if $\mathcal{P_{R}}$ has a width $\gtrsim 10^{-1}$, while below that the leading order of $y$ is much smaller than that of $x$. Therefore, we consider the region $x>\sqrt{6}/2$ first and regard $y$ as a perturbation in $F(x,y)$:
$\Omega_{\mathrm{GW}}^{x>\frac{\sqrt{6}}{2}}(k)$ $=$ $\frac{3}{4}\int_{\frac{\sqrt{6}}{2}}^{+\infty}\mathrm{d}x~{}x^{-16}\left(1-2x^{2}\right)^{2}\left(x^{2}-3\right)^{4}\left[\pi^{2}+\left(\frac{2x^{2}}{x^{2}-3}-\ln\left(\frac{2}{3}x^{2}-1\right)\right)^{2}\right]\mathcal{P}_{\mathcal{R}}^{2}(q)Q(x,q),$ (29)
[…]
$\approx$ $\int_{-\frac{1}{\sqrt{2}}}^{\frac{1}{\sqrt{2}}}\mathrm{d}y\left(1-2y^{2}\right)^{2}(1+4~{}\frac{y^{2}}{x^{2}})\mathcal{P_{R}}\left[q\left(1+\frac{y}{x}\right)\right]\mathcal{P_{R}}\left[q\left(1-\frac{y}{x}\right)\right]\mathcal{P}_{\mathcal{R}}^{-2}(q).$
$Q(x,q)$ refers to the contribution from the integration on $y$ relative to the $x=0$ slice. In the last step, the approximation is performed in the sense of integration, so that the singularity of $y^{2}$ order near $x=\sqrt{3}$ has been neglected, while it is guaranteed at $x=\sqrt{6}/2$ and $x\rightarrow+\infty$. The $x^{-2}$ dependence is selected for further convenience.
Similarly, we can expand the $\mathcal{P_{R}}$ factors in Eq. (30) up to $y^{2}/x^{2}$ order but a positive integrand should be guaranteed. Therefore, we replace the $x$-dependent factors in Eq. (30) with an exponential factor and mark the integral as $\bar{Q}(x,q)$:
$\bar{Q}(x,q)$ $=$ $\int_{-\frac{1}{\sqrt{2}}}^{\frac{1}{\sqrt{2}}}\mathrm{d}y\left(1-2y^{2}\right)^{2}\exp{\frac{4-\gamma(q)}{x^{2}}y^{2}}=\frac{\sqrt{\pi}\left(c^{2}+2c+3\right)\text{erfi}\left(\sqrt{c/2}\right)}{c^{5/2}}-\frac{\sqrt{2}(c+3)e^{c/2}}{c^{2}},$ (31)
@@ III.2 Analytical approach to SIGW
$c(x,q)$ $=$ $\frac{4-\gamma(q)}{x^{2}},\quad\gamma(q)=\alpha\beta\frac{\beta+\left(\alpha^{2}+2\alpha\beta+\alpha+(\beta-1)\beta\right)q^{\alpha+\beta}-\alpha q^{2(\alpha+\beta)}}{\left(\beta+\alpha q^{\alpha+\beta}\right)^{2}},$ (32)
where $\text{erfi}(z)=-i\text{erf}(iz)$ denotes the imaginary error function. $\gamma(q)$ denotes a constant-order transition with the asymptotic behavior that $\gamma(0)=\alpha$, $\gamma(1)=\alpha\beta$ and $\gamma(+\infty)=-\beta$, which are related to the IR tail, peak and the UV tail of the power spectrum respectively. Looking back to Eq. (29), it remains difficult if we use Eq. (31) directly, but it becomes simpler if we focus on the specific values of $\gamma(q)$ mentioned above. Also, it is helpful to take into account the approximation derived in Eq. (20) here, which also provides a segmentation to perform further approximation on $\bar{Q}(x,q)$.
In the case of $\alpha\beta\leq 6$, which means that the width of the peak on $y$ direction $(\alpha\beta\kappa^{2})^{-1/2}$ have not entered that of $(1-2y^{2})^{2}$ before the peak leaving this region, we take the expansion of Eq. (31) at $c=0$ to the first order:
$\bar{Q}(x,q)\approx\frac{8\sqrt{2}}{15}\left(1+\frac{1}{14}\frac{4-\gamma(q)}{x^{2}}\right).$ (33)
We simply let $\gamma(q)=\alpha\beta$ for $\kappa_{-}<q<\kappa_{+}$ as a near-peak case. Considering that the IR and UV part of $\mathcal{P_{R}}$ may also take over the leading order of $\Omega_{\mathrm{GW}}(k)$, we connect the asymptotic behavior of $\gamma(q)$ and then $\bar{Q}(x,q)$ can be approximated for $\alpha\beta\leq 6$:
$Q_{c}(x,\kappa)=\begin{cases}a_{0}+a_{-2}x^{-2}&\text{for}~{}x<\sqrt{2}\kappa_{-}/\kappa,\\ b_{0}+b_{-2}x^{-2}&\text{for}~{}\sqrt{2}\kappa_{-}/\kappa\leq x\leq\sqrt{2}\kappa_{+}/\kappa,\\ c_{0}+c_{-2}x^{-2}+c_{-4}x^{-4}&\text{for}~{}x>\sqrt{2}\kappa_{+}/\kappa,\end{cases}~{}~{}$ (34)
$a_{-2}=\frac{Q_{c}(\sqrt{2}\kappa_{-}/\kappa,\kappa)-\bar{Q}(\sqrt{6}/2,\sqrt{3}\kappa/2)}{(\sqrt{2}\kappa_{-}/\kappa)^{-2}-2/3},~{}a_{0}=Q_{c}(\sqrt{2}\kappa_{-}/\kappa,\kappa)-(\sqrt{2}\kappa_{-}/\kappa)^{-2}a_{-2},$
$b_{0}=c_{0}=\frac{8\sqrt{2}}{15},~{}b_{-2}=\frac{b_{0}}{14}(4-\alpha\beta),~{}c_{-2}=\frac{c_{0}}{14}(4+\beta),~{}c_{-4}=-\frac{c_{0}}{14}\beta(1+\alpha)\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{2}.$
Here we use $\kappa$ as a variable instead of $q$. The higher order $x^{-4}$ is introduced to guarantee the continuity of $Q_{c}(x,\kappa)$ at $x=\sqrt{2}\kappa_{+}/\kappa$, which will further guarantee the continuity of the derivative of $\Omega_{\mathrm{GW}}(k)$. As for the $x<\sqrt{2}\kappa_{-}/\kappa$ region, it is the first priority to ensure the accuracy at $x=\sqrt{6}/2$, and the asymptotic behavior on $\kappa\rightarrow 0$ turn into Eq. (33) naturally with $\gamma(q)=\alpha$.
@@
For the other case $\alpha\beta>6$ where the approximation on $c(x,q)\ll 1$ breaks down, we focus on the value of $\bar{Q}(x,q)$ at $q=1$ first:
$\bar{Q}(x,1)$$\approx$ $\left(\left(\frac{\sqrt{\pi}x}{\sqrt{\alpha\beta-4}}\right)^{-2}+\left(\frac{8\sqrt{2}}{15}\right)^{-2}\right)^{-1/2}.$ (35)
@@
We have connected the leading order of $x\ll\sqrt{\alpha\beta-4}$ and $x\gg\sqrt{\alpha\beta-4}$, with the relative error $<0.05$ compared with Eq. (31). We use the expansion of Eq. (35) at $x=\sqrt{2}/\kappa$ as a near-peak approximation for $\kappa_{-}<q<\kappa_{+}$:
@@
$Q_{c}(x,\kappa)=\left(\frac{1}{\kappa}\sqrt{\frac{2\pi}{(\alpha\beta-4)(\Delta_{\kappa}+1)}}\right)\frac{\Delta_{\kappa}+\kappa x/\sqrt{2}}{\Delta_{\kappa}+1},\quad\Delta_{\kappa}=\frac{225}{128}\frac{2\pi}{\kappa^{2}(\alpha\beta-4)}.~{}$ (36)
@@
The feature of Eq. (36) depend on the value of $\Delta_{\kappa}$. The value of the second factor of $Q_{c}(x,\kappa)$ falls between $\kappa_{-}$ and $\kappa_{+}$ and simply reduces to 1 when $x=\sqrt{2}/\kappa$ or $\Delta_{\kappa}\gg 1$, while the overall coefficient gives a transition from $\kappa^{-1}$ to constant near $\Delta_{\kappa}=1$, and the constant limit returns to that in Eq. (33). Similarly, we connect the asymptotic behavior to the near-peak approximation and the result for $\alpha\beta>6$ can be written as:
$Q_{c}(x,\kappa)=\begin{cases}a_{0}+a_{-2}x^{-2}&\text{for}~{}x<\sqrt{2}\kappa_{-}/\kappa,\\ b_{0}+b_{1}x&\text{for}~{}\sqrt{2}\kappa_{-}/\kappa\leq x\leq\sqrt{2}\kappa_{+}/\kappa,\\ c_{0}+c_{-2}x^{-2}+c_{-4}x^{-4}&\text{for}~{}x>\sqrt{2}\kappa_{+}/\kappa,\end{cases}$ (37)
$c_{-4}=\left(Q_{c}(\sqrt{2}\kappa_{+}/\kappa,\kappa)-c_{0}-c_{-2}\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{-2}\right)\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{4}.$
$a_{0}$, $a_{-2}$, $c_{0}$ and $c_{-2}$ have the same form with that in Eq. (34), while $b_{0}$ and $b_{1}$ can be obtained from Eq. (36).
@@
For both $\alpha\beta\leq 6$ and $\alpha\beta>6$, $Q(x,q)$ has been reduced to the sum of a group of $x^{\rho_{i}}$ terms and the coefficients have been expressed as $a_{\rho_{i}}$, $b_{\rho_{i}}$ and $c_{\rho_{i}}$ which is related to IR, near-peak and UV region of the power spectrum respectively. The contribution to $\Omega_{\mathrm{GW}}$ can be written down directly by combining Eqs. (20) (29) (34) and (37), while the parameters are also relevant to Eq. (31):
@@
$\Omega_{\mathrm{GW}}^{q<\kappa_{-}}(k)$$\approx$ $A^{2}(1+\lambda)^{2}\sum_{\rho_{i}}\sum_{j=0}^{2}a_{\rho_{i}}C_{2}^{j}d_{1}^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{2\alpha+j(\alpha+\beta)}\Omega(\rho_{i}+2\alpha+j(\alpha+\beta),x)\Big{|}_{\sqrt{6}/2}^{R_{1}},$ (38b)
$\Omega_{\mathrm{GW}}^{\kappa_{-}<q<1}(k)$$\approx$ $A^{2}\sum_{\rho_{i}}b_{\rho_{i}}\left(\Omega(\rho_{i},x)+d_{2}\sum_{j=0}^{2}C_{2}^{j}(-1)^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{j\beta}\Omega(\rho_{i}+j\beta,x)\right)\Bigg{|}_{R_{1}}^{R_{2}},$ (38c)
$\Omega_{\mathrm{GW}}^{1<q<\kappa_{+}}(k)$$\approx$ $A^{2}\sum_{\rho_{i}}b_{\rho_{i}}\left(\Omega(\rho_{i},x)+d_{3}\sum_{j=0}^{2}C_{2}^{j}(-1)^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{-j\alpha}\Omega(\rho_{i}-j\alpha,x)\right)\Bigg{|}_{R_{2}}^{R_{3}},$ (38d)
$\Omega_{\mathrm{GW}}^{q>\kappa_{+}}(k)$$\approx$ $A^{2}(1+\lambda^{-1})^{2}\sum_{\rho_{i}}\sum_{j=0}^{2}c_{\rho_{i}}C_{2}^{j}d_{4}^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{-2\beta-j(\alpha+\beta)}\Omega(\rho_{i}-2\beta-j(\alpha+\beta),x)\Big{|}_{R_{3}}^{+\infty}.$ (38e)
Here, $C_{m}^{n}$ represents the binomial coefficients. All other coefficients are listed in Appendix B for convenience, and $R_{1},~{}R_{2}$ and $R_{3}$ are defined as
@@
$\Omega(\rho,x)$ denotes the integration of $F(x,0)x^{\rho}$, which can be expressed as:
$\Omega(\rho,x)=$ $\frac{3\pi^{2}}{4}\sum_{m=0}^{4}\sum_{n=0}^{2}C_{4}^{m}C_{2}^{n}(-3)^{4-m}(-2)^{n}I(-16+2m+2n+\rho,0,x)$

### v2

@@ Appendix A Features of broken power-law
$\lambda=\alpha/\beta$ and $K(\kappa)=\kappa^{\alpha+\beta}\lambda$ are introduced for simplicity, and $\delta$ is the relative error of the power expansion up to $K^{n}$ term. We assume that the power-law feature is broken when the leading order of the asymptotic expansion reaches the peak value, where $\kappa$ and $\mathcal{P_{R}}(\kappa)$ satisfy that
@@
$\mathcal{P_{R}}(\kappa_{-})>\frac{A}{2},~{}\mathcal{P_{R}}(\kappa_{+})>\frac{A}{2},~{}\mathcal{P_{R}}(\kappa_{+})\mathcal{P_{R}}(\kappa_{-})>\frac{A^{2}}{2}.$
Here $\kappa_{-}$ and $\kappa_{+}$ are the critical values of IR and UV expansion respectively, and $\kappa_{0}$ is the boundary of the convergence domain of Eq. (47). The relation indicates that even if we use the power-law tail directly and replace the transition part $\kappa_{-}<\kappa<\kappa_{+}$ of the power spectrum by peak value, the relative error is smaller than 1. Moreover, the relative error approaches 1 on one side only if the power index is much larger than that on the other side, i.e., $|\ln\lambda|\gg 1$, and decays much more rapidly than the other side.
In order to elevate the approximation to a higher accuracy, we introduce a variable transformation that maintains $\mathcal{P_{R}}$:
$\mathcal{P_{R}}(\alpha,\beta;\kappa)=\mathcal{P_{R}}(\lambda,1;\kappa^{\beta})=\mathcal{P_{R}}(\lambda^{-1},1;\kappa^{-\alpha}).$ (49)
Therefore, once the approximation of $\mathcal{P_{R}}(\lambda,1,\kappa^{\beta})$ on $10<\kappa<$ 1 and $\lambda>0$ is finished, it can be extended to the complete parameter space $\alpha,\beta,\kappa>0$ with the accuracy preserved. The selection is based on that $\kappa_{-}^{\beta}=(1+\lambda)^{-1/\lambda}>e^{-1}$, which make it convenient to extend the power-law tail to the peak with a quadratic polynomial. On the region $\kappa<\kappa_{-}$, the deviation of the power-law tail from $\mathcal{P_{R}}$ is roughly to the order $(2\lambda+1)$, and consequently a general $\mathcal{P_{R}}(\kappa)$ can be approximated by:
@@
Additionally, $\mathcal{P_{R}}^{2}(\kappa)$ will be used in the calculation of $\Omega_{\mathrm{GW}}$, since it is associated with the $p=|\bm{k}-\bm{p}|$ slice in Eq. (5). We use the similar approximation:
@@
We have used that $\kappa_{-}(\lambda,1)=\kappa_{-}^{\beta}(\alpha,\beta)$. Another integral will also be used:
$\int_{0}^{+\infty}\mathcal{P_{R}}(t)\mathcal{P_{R}}(t^{-1})\mathrm{d}\ln t=\int_{0}^{+\infty}\frac{\alpha+\beta}{(\alpha+\beta s)(\alpha s+\beta)}\mathrm{d}s=\frac{2}{\sqrt{\alpha\beta}}\frac{\ln\lambda}{\lambda^{1/2}-\lambda^{-1/2}}.$ (54)
The second factor of the result is not sensitive to $\lambda$, and the value in the limit of $\lambda\to 1$ is simply 1.
@@
Appendix B Contributions from region $x>\sqrt{6}/2$
We begin from Eqs. (18) and (19). If the enhancement in the power spectrum is relatively slow, we can expand the integrand in Eq. (19) up to $y^{2}/x^{2}$ order, and hence $I_{y}(x,q)$, where $q=\kappa x/\sqrt{2}$, can be derived analytically:
@@ Appendix B Contributions from region $x>\sqrt{6}/2$
$I_{y}(x,q)\approx\frac{8\sqrt{2}}{15}\left(1+\frac{1}{14}\frac{4-\gamma(q)}{x^{2}}\right),\quad\gamma(q)=\alpha\beta\frac{\beta+\left(\alpha^{2}+2\alpha\beta+\alpha+(\beta-1)\beta\right)q^{\alpha+\beta}-\alpha q^{2(\alpha+\beta)}}{\left(\beta+\alpha q^{\alpha+\beta}\right)^{2}}.$ (60)
The above approximation is valid if $\alpha\beta\leq 6$, which means that the width of the power spectrum $(\alpha\beta\kappa^{2})^{-1/2}$ exceeds that of $(1-2y^{2})^{2}$ before the peak leaves this region. $\gamma(q)$ denotes a constant-order transition and we will use the asymptotic behavior that $\gamma(0)=\alpha$, $\gamma(1)=\alpha\beta$ and $\gamma(+\infty)=-\beta$, which are related to the IR tail, peak and the UV tail of the power spectrum respectively.
According to the segmentation in Eq. (15), we use Eq. (60) directly with $\gamma(q)=\gamma(1)$ for the near-peak case $\kappa_{-}\leq q\leq\kappa_{+}$, and denote as $I_{y,p}(x)$:
$I_{y,p}(x)=\frac{8\sqrt{2}}{15}\left(1+\frac{1}{14}\frac{4-\alpha\beta}{x^{2}}\right).$ (61)
Then we connect the behavior of $\gamma(q)$ at $q=\sqrt{3}\kappa/2$ and $q\to+\infty$ to the near-peak approximation above. The selection is based on the boundary of the integration on $x$. In this case, $I_{y}(x,q)$ can be approximated to:
$I_{y}(x,q)\approx\begin{cases}a_{0}+a_{-2}x^{-2}&\text{for}~{}x<\sqrt{2}\kappa_{-}/\kappa,\\ b_{0}+b_{-2}x^{-2}&\text{for}~{}\sqrt{2}\kappa_{-}/\kappa\leq x\leq\sqrt{2}\kappa_{+}/\kappa,\\ c_{0}+c_{-2}x^{-2}+c_{-4}x^{-4}&\text{for}~{}x>\sqrt{2}\kappa_{+}/\kappa,\end{cases}~{}~{}$ (62)
$a_{-2}=\frac{I_{y,p}(\sqrt{2}\kappa_{-}/\kappa)-I_{y}(\sqrt{6}/2,\sqrt{3}\kappa/2)}{(\sqrt{2}\kappa_{-}/\kappa)^{-2}-2/3},~{}a_{0}=I_{y,p}\left(\sqrt{2}\kappa_{-}/\kappa\right)-a_{-2}\left(\frac{\sqrt{2}\kappa_{-}}{\kappa}\right)^{-2},$
$b_{0}=c_{0}=\frac{8\sqrt{2}}{15},~{}b_{-2}=b_{0}\frac{1}{14}(4-\alpha\beta),~{}c_{-2}=c_{0}\frac{1}{14}(4+\beta),~{}c_{-4}=-\frac{c_{0}}{14}\beta(1+\alpha)\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{2}.$
The intervals correspond to $q<\kappa_{-}$, $\kappa_{-}\leq q\leq\kappa_{+}$, $q>\kappa_{+}$ respectively. The coefficients $a_{\rho}$, $b_{\rho}$ and $c_{\rho}$, where $\rho$ denote the corresponding power index of $x$, are relevant to $\kappa$ and the shape of power spectrum. In the region $x<\sqrt{2}\kappa_{-}/\kappa$, we connect the value of $I_{y}$ at $x=\sqrt{6}/2$ to the near-peak case in Eq. (61), and the asymptotic behavior on $\kappa\rightarrow 0$ will turn into Eq. (60) naturally with $\gamma(q)\to\alpha$. Here the value of $I_{y}(\sqrt{6}/2,\sqrt{3}\kappa/2)$ will be given in Eq. (76). In the region $x>\sqrt{2}\kappa_{+}/\kappa$, we use Eq. (60) and introduce an extra $x^{-4}$ term to guarantee the continuity at $x=\sqrt{2}\kappa_{+}/\kappa$.
@@
For the other case $\alpha\beta>6$, we focus on the peak value $I_{y}(x,1)$ first:
$I_{y}(x,1)$$\approx$ $\left(\left(\frac{2x}{\sqrt{\alpha\beta-4}}\right)^{-2}+\left(\frac{8\sqrt{2}}{15}\right)^{-2}\right)^{-1/2}.$ (63)
@@
We have connected the leading order of $2x\ll\sqrt{\alpha\beta-4}$ and $2x\gg\sqrt{\alpha\beta-4}$, with the relative error $<0.1$ compared with the normal approximation in Eq. (76). We use the expansion of Eq. (63) at $x=\sqrt{2}/\kappa$ as a near-peak approximation for $\kappa_{-}<q<\kappa_{+}$:
@@
$I_{y,p}(x)=\left(\frac{2}{\kappa}\sqrt{\frac{2}{(\alpha\beta-4)(\Delta_{\kappa}+1)}}\right)\frac{\Delta_{\kappa}+\kappa x/\sqrt{2}}{\Delta_{\kappa}+1},\quad\Delta_{\kappa}=\frac{225}{16}\frac{1}{\kappa^{2}(\alpha\beta-4)}.~{}$ (64)
@@
The feature of Eq. (64) depends on the value of $\Delta_{\kappa}$. The value of the second factor of $I_{y,p}$ falls between $\kappa_{-}$ and $\kappa_{+}$ and simply reduces to 1 when $x=\sqrt{2}/\kappa$ or $\Delta_{\kappa}\gg 1$, while the overall coefficient gives a transition from $\kappa^{-1}$ to constant near $\Delta_{\kappa}=1$, and the constant limit returns to that in Eq. (60).
Similarly, we connect the asymptotic behavior to the near-peak approximation and the result for $\alpha\beta>6$ can be written as:
$I_{y}(x,q)\approx\begin{cases}a_{0}+a_{-2}x^{-2}&\text{for}~{}x<\sqrt{2}\kappa_{-}/\kappa,\\ b_{0}+b_{1}x&\text{for}~{}\sqrt{2}\kappa_{-}/\kappa\leq x\leq\sqrt{2}\kappa_{+}/\kappa,\\ c_{0}+c_{-2}x^{-2}+c_{-4}x^{-4}&\text{for}~{}x>\sqrt{2}\kappa_{+}/\kappa,\end{cases}$ (65)
$b_{0}=\frac{2}{\kappa}\sqrt{\frac{2}{(\alpha\beta-4)(\Delta_{\kappa}+1)}}\frac{\Delta_{\kappa}}{\Delta_{\kappa}+1},\quad b_{1}=2\sqrt{\frac{1}{(\alpha\beta-4)(\Delta_{\kappa}+1)}}\frac{1}{\Delta_{\kappa}+1}$
$c_{-4}=\left(I_{y,p}(\sqrt{2}\kappa_{+}/\kappa)-c_{0}-c_{-2}\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{-2}\right)\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{4}.$
$a_{0}$, $a_{-2}$, $c_{0}$ and $c_{-2}$ have the same form with that in Eq. (62).
@@
For both $\alpha\beta\leq 6$ and $\alpha\beta>6$, $I_{y}(x,q)$ has been reduced to the sum of a group of $x^{\rho}$ terms and the coefficients have been expressed as $a_{\rho}$, $b_{\rho}$ and $c_{\rho}$ which is related to IR, near-peak and UV region of the power spectrum respectively. Taking into account the approximation in Eq. (15), the contribution to $\Omega_{\mathrm{GW}}$ from this region can be written down directly by combining Eqs. (15) (17) (18) (62) and (65):
@@
$\Omega_{\mathrm{GW}}^{q<\kappa_{-}}(k)$$\approx$ $A^{2}(1+\lambda)^{2}\sum_{\rho}\sum_{j=0}^{2}a_{\rho}C_{2}^{j}d_{1}^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{2\alpha+j(\alpha+\beta)}\Omega(\rho+2\alpha+j(\alpha+\beta),x)\Big{|}_{\sqrt{6}/2}^{R_{1}},$ (66b)
$\Omega_{\mathrm{GW}}^{\kappa_{-}<q<1}(k)$$\approx$ $A^{2}\sum_{\rho}b_{\rho}\left(\Omega(\rho,x)+d_{2}\sum_{j=0}^{2}C_{2}^{j}(-1)^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{j\beta}\Omega(\rho+j\beta,x)\right)\Bigg{|}_{R_{1}}^{R_{2}},$ (66c)
$\Omega_{\mathrm{GW}}^{1<q<\kappa_{+}}(k)$$\approx$ $A^{2}\sum_{\rho}b_{\rho}\left(\Omega(\rho,x)+d_{3}\sum_{j=0}^{2}C_{2}^{j}(-1)^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{-j\alpha}\Omega(\rho-j\alpha,x)\right)\Bigg{|}_{R_{2}}^{R_{3}},$ (66d)
$\Omega_{\mathrm{GW}}^{q>\kappa_{+}}(k)$$\approx$ $A^{2}(1+\lambda^{-1})^{2}\sum_{\rho}\sum_{j=0}^{2}c_{\rho}C_{2}^{j}d_{4}^{j}\left(\frac{\kappa}{\sqrt{2}}\right)^{-2\beta-j(\alpha+\beta)}\Omega(\rho-2\beta-j(\alpha+\beta),x)\Big{|}_{R_{3}}^{+\infty}.$ (66e)
All the defined coefficients are listed below for convenience. Here, $C_{m}^{n}$ represents the binomial coefficients, $\lambda=\alpha/\beta$, and $R_{1},~{}R_{2}$ and $R_{3}$ are defined as
@@
$\Omega(\rho,x)$ denotes the integration of $F(x,0)x^{\rho}$. Actually, the calculation can be done directly and the analytical results are listed in Appendix C.
For the coefficients relevant to the segment of the scalar power spectrum which is derived in Appendix A:
$\mathcal{P_{R}}(q)=A\frac{\alpha+\beta}{\beta q^{-\alpha}+\alpha q^{\beta}},~{}\kappa_{-}=\left(1+\lambda\right)^{-1/\alpha},~{}\kappa_{+}=\left(1+\lambda^{-1}\right)^{1/\beta},$ (68)
Then for the coefficients relevant to the integration on $y$ direction, i.e. $I_{y}(x,q)$, the coefficients in each region is defined in Eq. (62) and (65). For the contribution from the region $\kappa_{-}\leq q\leq\kappa_{+}$, the relevant coefficients in Eq. (66c) and Eq. (66d) are that
$(\rho,b_{\rho})\in\begin{cases}\left\{\left(0,\dfrac{8\sqrt{2}}{15}\right),\left(-2,\dfrac{b_{0}}{14}(4-\alpha\beta)\right)\right\},&\text{for}~{}\alpha\beta\leq 6;\\ \\ \left\{\left(0,\sqrt{\dfrac{8}{(\alpha\beta-4)}}\dfrac{\kappa^{-1}\Delta_{\kappa}}{(\Delta_{\kappa}+1)^{3/2}}\right),\left(1,\sqrt{\dfrac{4}{(\alpha\beta-4)}}\dfrac{1}{(\Delta_{\kappa}+1)^{3/2}}\right)\right\},&\text{for}~{}\alpha\beta>6.\end{cases}$ (70)
$\Delta_{\kappa}$ is defined as
$\Delta_{\kappa}=\frac{225}{16}\frac{1}{\kappa^{2}(\alpha\beta-4)}.$ (71)
$(\rho,c_{\rho})\in\left\{\left(0,\frac{8\sqrt{2}}{15}\right),\left(-2,\frac{c_{0}}{14}(4+\beta)\right),\left(-4,\left[I_{y,p}\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)-c_{0}-c_{-2}\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{-2}\right]\left(\frac{\sqrt{2}\kappa_{+}}{\kappa}\right)^{4}\right)\right\},$ (72)
where $I_{y,p}(x)$ is the near-peak approximation, which is expressed as
$I_{y,p}\left(\frac{\sqrt{2}\kappa_{\pm}}{\kappa}\right)=\begin{cases}\dfrac{8\sqrt{2}}{15}\left(1+\dfrac{4-\alpha\beta}{14}\dfrac{\kappa^{2}}{2\kappa_{\pm}^{2}}\right)&\text{for}~{}\alpha\beta\leq 6,\\ \\ \dfrac{2}{\kappa}\sqrt{\dfrac{2}{(\alpha\beta-4)(\Delta_{\kappa}+1)}}\dfrac{\Delta_{\kappa}+\kappa_{\pm}}{\Delta_{\kappa}+1}&\text{for}~{}\alpha\beta>6,\end{cases}$ (73)
It is also used that for $\alpha\beta\leq 6$, $c_{-4}$ reduces to that
$(\rho,a_{\rho})\in\left\{\left(0,I_{y,p}\left(\frac{\sqrt{2}\kappa_{-}}{\kappa}\right)-\left(\frac{\sqrt{2}\kappa_{-}}{\kappa}\right)^{-2}a_{-2}\right),\left(-2,\frac{I_{y,p}(\sqrt{2}\kappa_{-}/\kappa)-I_{y}(\sqrt{6}/2,\sqrt{3}\kappa/2)}{(\sqrt{2}\kappa_{-}/\kappa)^{-2}-2/3}\right)\right\}.$ (75)
The value of $I_{y}(\sqrt{6}/2,\sqrt{3}\kappa/2)$ is quite important here. Beginning from Eq. (19), we treat the peak of power spectrum on $y$ direction as normal and the result is that
$I_{y}(x,q)\approx\frac{\sqrt{\pi}\left(c^{2}+2c+3\right)\text{erfi}\left(\sqrt{c/2}\right)}{c^{5/2}}-\frac{\sqrt{2}(c+3)e^{c/2}}{c^{2}},$ (76)
where $\text{erfi}(z)=-i\text{erf}(iz)$ denotes the imaginary error function.
Appendix C Analytical expressions of $\Omega(\rho,x)$
In this section, we give the analytical expression of $\Omega(\rho,x)$, which has been used in the calculation on $\Omega_{\mathrm{GW}}(k)$ in Eqs. (66a)-(66e), Appendix B. $\Omega(\rho,x)$ is defined as
$\Omega(\rho,x)=$ $\int\mathrm{d}x~{}x^{\rho}F(x,0)$ (78)
$=$ $\frac{3\pi^{2}}{4}\sum_{m=0}^{4}\sum_{n=0}^{2}C_{4}^{m}C_{2}^{n}(-3)^{4-m}(-2)^{n}I(-16+2m+2n+\rho,0,x)$